OpenStudy (anonymous):

Let , x^2+2x, x is > or = to -1. Find its inverse and state its domain and range using interval notation. If required, use fractions, NOT decimals. I did quadratic formula and got: My answer: (-2+-sqrt(4+4x))/2 Domain of : (-oo,-1] but this answer is wrong, what did I do wrong?

OpenStudy (anonymous):

I put [-1,oo) for range and that was correct...

OpenStudy (anonymous):

Yeah, the range is \([-1,\infty)\).

OpenStudy (anonymous):

\[\left( -2\pm \sqrt(2^2-4(1)(-1)\right)/2\] This is my answer if thats clearer

OpenStudy (anonymous):

hello anwar! mistake is the \[\pm\] this a function so only one output for each input.

OpenStudy (anonymous):

To find the inverse, set \(x=y^2+2y\), and try to solve for y. You will have \(y^2+2y-x=0\), by using the quadratic formula, we have: \[y={-2\pm \sqrt{4+4x} \over 2}=-1\pm \sqrt{1+x}\]. But since the range is \((-1,\infty)\); we will take only \(+\sqrt{1+x}\).

OpenStudy (anonymous):

Hello satellite!

OpenStudy (anonymous):

hello. nice latex.

OpenStudy (anonymous):

So, the inverse is \(\sqrt{1+x}-1\). Try to find the domain. And thank ssatellite :)

OpenStudy (anonymous):

ok. S owhy would the answer for the domain not be (-oo,-1]?

OpenStudy (anonymous):

The domain of the inverse has to be the same as the range of the original function. So, what do you think is the range of the original function, in the given domain \((-1,\infty)\)?

OpenStudy (anonymous):

\([-1,\infty)\)*

OpenStudy (anonymous):

why is \[\pm\] a mistake?

OpenStudy (anonymous):

A function can't have a ±, because it would fail the vertical line test then.

OpenStudy (anonymous):

The domain of the inverse, by the way, is \([-1,\infty)\).

OpenStudy (anonymous):

oh, but I thought I'm to use quadratic formula to find the inverse of f(x)

OpenStudy (anonymous):

in this situation

OpenStudy (anonymous):

let me butt in for a second. think about \[f(x)=x^2\] if you write \[x=y^2\] and solve for y you get \[y=\pm\sqrt{x}\] but this is not a function. and of course \[f(x)=x^2\] is not a one to one function. but if you restrict the domain of \[f(x)\] to x>0, then the inverse has to have range y > 0 and so you just write \[f^{-1}(x)=\sqrt{x}\] a perfectly good function.

OpenStudy (anonymous):

I understand that there is only one output for one input, so how am i supposed to do that?

OpenStudy (anonymous):

But, in the original function the domain is \([-1,\infty)\), and therefore the range of the inverse has to be \([-1,\infty)\). But if we take \(-\sqrt{x+1}\), that would lead to values that are not part of the range.

OpenStudy (anonymous):

how do I express that?

OpenStudy (anonymous):

It's already expressed :D

OpenStudy (anonymous):

so, my question i guess is what is the range? [-1,oo)?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Im confused. If it leads to wrong values, then why does it work?

OpenStudy (anonymous):

Im confused. If it leads to wrong values, then why does it work?

OpenStudy (anonymous):

That would be only If we take the minus the square as part of the solution, but we didn't. So, we're in the safe side.

OpenStudy (anonymous):

OOOOO! I gotcha, thank you :)

OpenStudy (anonymous):

can I ask you different question, about absolute value? x=|y+3|

OpenStudy (anonymous):

trying to figure out inverse of that?